# Manipulation Priors Learned From Egocentric Videos

### v1

@@ 4.3 Evaluation in Real-World Settings
Figure 5: Real-World Sequences of the Evaluated Tasks. Example rollouts of MAPLE. From top to bottom: ‘Grab the plush animal (plush)’, ‘Place the pan’ (pan) and ‘paint the canvas’ (brush) tasks using the Mimic hand and the Franka Emika Panda manipulator.
Task Designs. Our real-world evaluation features two tasks derived from those used in simulation, as well as a novel task: (1) Plush: The robot picks up a plush object and places it on a tray, assessing the robustness to objects with a unique shape unseen during the encoder’s training. (2) Pan: The robot grasps the handle of a pan, lifts it, and places it onto an induction stove. After placement, the robot releases the handle. (3) Brush: The robot picks up a brush, moves it to a target frame, and performs a brushing motion on its surface before dropping the brush on a table. This is the most challenging task, testing both grasping precision and fine manipulation dexterity. Additional task descriptions and data collection details are in Supp. Mat. Sec. S1.
@@
Hardware Setup. Our robotic setup comprises a Mimic robotic hand [36, 60] mounted on a Franka Emika Panda robotic arm [18], along with two external cameras. Demonstrations are collected using Rokoko Smartgloves and the Coil Pro motion capture system [49], which provide proprioceptive location data from the demonstrator’s fingers and wrist. The external cameras record visual data.
@@
Experiment Protocol. After training on egocentric videos, we freeze the encoder to ensure stable feature extraction and use the extracted features for manipulation policy training. For our real-world experiments, we employ the Diffusion Policy framework [5] to predict robot actions and use the Diffusion Transformer (DiT) [12] as the model architecture. Visual inputs are processed by our trained encoders, and their embeddings are combined with proprioceptive data. In particular, the end-effector pose and finger joint angles are transformed into tokens and concatenated with the image tokens before being passed into the DiT. We use DDiM [54] as our noise scheduler, with 100 steps during training and 8 steps at inference. Additional architectural and training details are provided in Supp. Mat. S1.2.
Results. For each model and task, we conducted 16 trials to evaluate success rates. Figure 6 summarizes the performance across all models. Our results indicate that our encoders consistently outperform alternative approaches. In particular, MAPLE-LN shows the highest overall performance, benefiting from its stability and strong generalizability across varying task conditions. Meanwhile, MAPLE-AP slightly outperforms MAPLE-LN on the most dexterous brush task. In contrast, other encoder-based methods exhibit behaviors that make them less suitable for our tasks: Using R3M [40] results in the policy overfitting and being less responsive to variations in initial conditions, leading to the lowest success rates. Using HRP [55] yields moderate performance on the pan task but induces abrupt hand and wrist movements, indicating challenges in accurate spatial reasoning. DINO [4] and VIP [31] offer competitive performance; however, they do not consistently match the stability and adaptability of MAPLE. VIP shows a slight edge over DINO on the brush task, but overall, both lag behind MAPLE in balancing generalizability and task-specific dexterity. VC-1 [32] presents unstable performance, performing well on the pan task yet struggling with the plush and brush tasks, seemingly insensitive to those tasks’ vision clues. Overall, our results demonstrate that MAPLE offers clear benefits in terms of both performance and adaptability not only in simulation but also in real-world manipulation tasks. Videos of our real-world experiments for all encoders are available in the Supp. Mat.
@@
Figure 6: Results on Real-World Environments. We evaluate different encoders and use the extracted features for imitation learning. Missing bars correspond to a zero success rate.
@@ 5 Conclusion
In this work, we study how to infuse dexterous manipulation priors into visual representations by training on diverse human videos, aiming to improve performance on downstream robotic tasks in human-centric environments. Common jaw grippers [8, 2, 1] limit interactions with a wide range of objects, making human-like hands more suitable for operating in human environments. We introduced MAPLE, a novel approach that learns manipulation-specific priors from egocentric videos. MAPLE trains an encoder to predict hand-object contact points and detailed 3D hand poses from a single image, producing features useful for downstream manipulation tasks. For the benefit of the community and a more rigorous evaluation, we further propose new dexterous simulation environments. Our strong results on both new tasks and existing benchmark tasks in simulation, as well as our superior real-world performance demonstrate promising manipulation prior learning and excellent generalization ability with MAPLE. Future directions of research include combining MAPLE with language conditioning and incorporating more diverse, possibly exocentric human manipulation datasets during its training.
@@ 6 Acknowledgements
This work was supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a03 on Alps, as well as the Swiss National Science Foundation Advanced Grant 216260: “Beyond Frozen Worlds: Capturing Functional 3D Digital Twins from the Real World”. The authors wish to express their gratitude to mimic robotics [36] for providing the robotic hand used in the real-world experiments.
@@
Figure S7: Initial Setups of All Real-World Evaluations. Initial setups for Plush Toy Pick and Place, Pan Pick and Place, and Frame Brush. Each method is evaluated for 16 episodes, all starting from the states shown in the figure.
In this section, we provide more details on the imitation learning evaluations introduced in Sec. 4.3. Each method is evaluated for 16 episodes for every task, with a predefined initial setup as shown in Fig. S7.
@@
Our system consists of a Mimic robotic hand [36] mounted on a Franka Emika Panda robotic arm [18]. Two external OAK-D cameras capture the scene, and the arm and hand are controlled at 10 Hz. The arm’s action is defined by a 7-dimensional end-effector pose ($x$, $y$, $z$, and a quaternion), and the hand’s action is represented by a 16-dimensional vector corresponding to its joint angles. All camera images are cropped and resized to 224×224 pixels. Our framework uses an observation window spanning two time steps, while control actions are executed in chunks spanning 10 time steps.
S1.2 Imitation Learning Details
@@
To accelerate training and enable larger batch sizes, we cache embeddings from the observation encoders directly within the dataset. This approach significantly speeds up training. Our DiT [12] backbone has a model dimension of 512, consists of 6 blocks, and employs 8 attention heads. The network is trained using the Adam optimizer with a cosine warmup learning rate schedule. All methods are evaluated after 35,000 training steps.
S1.3 Plush Pick and Place
@@
Demonstrations. 101 demonstrations are collected and used for training. 80% of them are used as training set, 20% as evaluation set. The starting condition is randomized. Demonstrations are collected by an experienced operator wearing Rokoko Smartgloves and the Coil Pro motion capture setup [49], which provides proprioceptive location data of the demonstrator’s fingers and wrist. Images from three cameras are collected and fed during policy rollout: side camera, front camera and wrist camera.
Evaluation. Each method is evaluated for 16 episodes, all starting from the same set of initial conditions as shown in Fig. S7.a. For each trial and task, the robot starts from the nominal pose. The episode is counted as successful only if the toy is grasped and fully placed in the tray. A second attempt is allowed if the policy misses the first grab and successfully grabs the toy without human intervention within the time limit of 30 seconds.
@@
S1.4 Pan Pick and Place
Demonstrations. 446 demonstrations are collected and used for training. 80% of them are used as training set, 20% as evaluation set. The starting condition is randomized. The demonstrations are collected with the same Rokoko solution as described before. Images from two cameras are collected and fed during policy rollout: side camera and front camera.
Evaluation. Each method is evaluated for 16 episodes, all starting from the same set of initial conditions as shown in Fig. S7.b. For each trial and task, the robot starts from the nominal pose. The episode is counted as successful only if the pan is grasped, lifted, placed on the induction stove and touches nothing else.
S1.5 Brush Frames
Demonstrations. 226 demonstrations are collected and used for training. 80% of them are used as training set, 20% as evaluation set. The starting condition is randomized. The demonstrations are collected using the same Rokoko solution as described above. Images from two cameras are collected and fed during the policy rollout: side camera and front camera.
Evaluation. Each method is evaluated for 16 episodes, all starting from the same set of initial conditions as shown in Fig. S7.c. For each trial and task, the robot starts from the nominal pose. The episode is counted as successful only if the brush is successfully grasped, touches the board, and is then dropped on the table.
S2 Demonstration Data and Benchmark Environments for Simulation
@@
S2.1 Hand Retargeting Algorithm
@@
The algorithm from [44] works by optimizing over the vector of the Adroit joint actuations, minimizing a Huber loss on the difference between pairs of an inter-fingertip or finger-to-wrist vector on the human hand and a corresponding vector measuring the same distance on the robotic hand. The hand pose is inferred by the aforementioned retargeting procedure, while the global position of the wrist joint is obtained using the Rokoko Coil Pro, an EMF-based tracking device. We use virtual PID controllers to drive each joint.
S2.2 Visualization of Demonstration Trajectories
@@
We show selected frames from our demonstration trajectories for the brush (Figure S8), the iron (Figure S9) and the pan (Figure S10) tasks for each camera view.
S3 Evaluation Details in Simulation
@@
To evaluate the suitability of a given visual encoder for downstream dexterous manipulation tasks, we use simulators in which a policy network consuming features from a visual encoder is trained to operate an observed robotic manipulator. As can also be expected from a real-world deployment, the policy network additionally receives proprioceptive features describing the current joint configurations of the robot as input. Position and orientation information related to the object(s) to be manipulated is only provided through visual cues, so as to increase realism and encourage the use of visual features to grasp and manipulate the object. The Franka Kitchen [17] and Metaworld [67] environments feature Franka Emika Panda [18] arms (7 DoF) with jaw grippers, exhibiting 7+2 DoF, while the DAPG and our dexterous environments feature the Adroit [24] dexterous hand with 30 DoF in total.
@@
The policy networks are trained to perform the respective tasks using demonstrations recorded by human participants, as described in Sec. 4.2. While training the policy networks, we periodically perform rollouts (i.e. simulation runs where the performance of the policy is assessed). The policies are trained for 20,000 training iterations. Rollouts are performed every 1,000 training iterations, and the highest success rate achieved over a given run is chosen for the computation of the final average success rates reported in the tables. 50 rollouts and 2 views are used during each evaluation for the experiments on the Franka Kitchen environments in Table 2. Note that the MetaWorld and DAPG environments, as well as our dexterous environments randomly set the orientation and position of the object(s) to manipulate upon each simulation restart to make the tasks more challenging and realistic.
@@
During the training of the policy, the checkpoint of the respective visual encoder is used as a frozen feature extractor, and only the parameters of the policy are trained. For our MAPLE-LN model, we use the checkpoint obtained after training 3,000 iterations on our Ego4D-based dataset. For MAPLE-AP, we use the checkpoint obtained after 2,000 iterations. These iteration counts were chosen to match the approximate contact prediction performance saturation point on a small validation dataset, manually constructed from prediction frames produced by our data extraction pipeline. In all simulation experiments, the policy is a 2-layer MLP with 256 units for each hidden layer. The MLP receives the visual features produced by the respective encoder from a rendered RGB view of the input scene. Examples of rendered simulator views are visualized in Figure S8, Figure S9, Figure S10. The MLP further receives a number $p$ of proprioceptive features encoding the current joint configuration of the jaw gripper resp. robotic hand. Specifically, we use $p=9$ for all Franka Kitchen environments and $p=4$ for all Metaworld environments. We use $p=24$ for the DAPG pen environment, as well as $p=30$ for the DAPG relocate, hammer and door environments. These values are chosen to match the setup in previous work [40]. We further set $p=30$ for all our dexterous environments. The number of proprioceptive features generally corresponds to the DoF of the robotic manipulator. At each simulation timestep, given these environment observations, the MLP networks output actuation vectors which determine the force applied at each actuator according to the conversion outlined in [38]. To learn from human demonstrations, the policy is provided observation inputs recorded during human teleoperation of the robot manipulator, and is trained to predict the actuation vector corresponding to the retargeted action that the human performed at that timestep. As is standard in literature [47, 40], we train a separate policy per environment.
S3.1 Environment Success Criteria
@@
Here, we provide detailed criteria that must be fulfilled for a task to count as successfully completed, for each of our dexterous environments. For distance measurements, MuJoCo site markers [38] attached to the robotic hand, manipulated object (brush, pan, iron) and target object (wooden board, induction stove, clothes) are used.
@@
  • Clothes Iron: The bottom of the clothes iron must slide along at least three of eight markers placed equidistantly along the right leg of the pants laid out on the table, while the palm of the hand is in close vicinity of the handle of the clothes iron.
S3.2 Limitations
Our work only considers a single embodiment in simulation and a single embodiment in the real-world experiments, making it difficult to estimate the impact of the particular robot on the performance achievable with our method. Further, the training data extracted using our pipeline exhibits noise (see subsection S4.1, subsection S4.2) that is likely harmful during training of the encoder.
S4 Training Supervision Extraction
@@
To obtain data for training our visual encoder, we start by processing every frame in the clips subset of the Ego4D dataset [15] using the VISOR-HOS [10] hand-object interaction segmenter, and retain frames where exactly one right hand is contacting an object with an HOS confidence at least $c\geq 0.9$, and no left hand is contacting an object with a similarly high confidence. Afterwards, we obtain frame-wise HaMeR [42] hand reconstructions for each contact frame. These hand reconstructions are used to initialize the prediction-frame–seeking contact point tracking with the thumb and index fingertips: for each contact frame, we first perform a binary erosion operation (12 iterations) on the object mask to remove the boundary, as motivated in Section 3.1. Then, we project the thumb and index fingertip points of the right hand in the contact frame to the eroded object mask. In case the ratio of the projected points’ Euclidean distance to their original distance falls outside the range between 0.3 and 1.7, the contact frame is not used. This helps eliminate objects with degenerate HOS masks. Otherwise, we initialize SAM-PT [48] with the projected fingertip points, as well as 10 other points sampled from the eroded object mask according to the query point sampling algorithm described in [48] and used to track the object. We continue tracking the points backward through the video until we encounter a frame where the HOS hand mask dilated using 75 binary dilation iterations (i.e., the expanded hand mask) no longer intersects the backtracked object mask. In case no such frame is encountered after 45 frames (i.e. 1.5s), the given contact frame is discarded. Otherwise, we use this frame as the prediction frame, and the backtracked points in that frame as training supervision for the contact loss. For the hand pose loss, we note that there are few constraints on the hand pose as it is moving in the direction of an object to grasp it. Hence, we always let our method reconstruct the hand pose at the contact frame, to promote learning the more informative grasping pose expected to be seen when the hand is already in contact with the object.
@@
Figure S12: Object Mask and Contact Points. The left column displays the prediction frames, while the right column shows the corresponding contact frames automatically extracted during our dataset acquisition step. The cyan cross mark represents the contact point of the thumb, and the pink cross mark indicates the contact point of the index finger.
S4.1 Error Analysis of Contact Supervision Extraction

### v2

@@ 4.3 Evaluation in Real-World Settings
Figure 5: Real World Sequences of the Evaluated Tasks. Example rollouts of MAPLE. From top to bottom: ‘Open the bottle cap’ (bottle), ‘Place the pan‘ (pan) and ‘Wash the dish‘ (sponge) tasks using the ORCA hand and the Franka Emika Panda manipulator.
Task Designs. Our real-world evaluation features one task derived from those used in simulation, as well as two novel tasks: (1) Pan: The robot grasps the handle of a pan, lifts it, and places it onto a burner. After placement, the robot releases the handle. This task features an object commonly found in the Ego4D dataset. (2) Bottle: The robot needs to open the cap of a bottle. This requires executing a twisting motion after a correct pinching grasp of the cap.
(3) Sponge: The robot picks up a sponge, moves it to an adjacent plate, sets it down to make contact with the plate’s surface, and circles the sponge around the plate while maintaining the grasp. Additional task descriptions and data collection details can be found in Supp. Mat.
@@
Hardware Setup. Our setup comprises an ORCA robotic hand [ORCA] mounted on a Franka Emika Panda arm [FrankaEmika], along with an external RGB camera and a wrist-mounted RGB camera located below the hand. Human demonstrations are collected using Rokoko Smartgloves and the Coil Pro motion capture system [Smartglove], which provide proprioceptive location data from the demonstrator’s fingers and wrist.
@@
Experiment Protocol. After training on egocentric videos, we freeze our encoder to ensure stable feature extraction and use the extracted features for manipulation policy training. We employ the Diffusion Policy framework [chi2023diffusion] to predict robot actions using the Diffusion Transformer (DiT) [dasari2024DiTpolicy]. Visual inputs are processed by our trained encoders, and their embeddings are combined with proprioceptive data. We use DDiM [Song2021ddim] as our noise scheduler, with 100 steps during training and 8 steps during inference. Further architectural and training details are in Supp. Mat.
@@
Figure 6: Results on Real-World Environments. We evaluate different encoders, using the extracted features for imitation learning. Missing bars indicate zero success rates.
Results. For each model and task, we conduct trials using 16 settings to evaluate success rates, each with a different position and/or rotation of the objects in the scene. Figure 6 summarizes the performance across all models. Our results indicate that our encoder consistently outperforms alternative approaches. MAPLE shows the highest overall performance, benefiting from its stability and strong generalizability across varying task conditions. In contrast, other encoder-based methods exhibit behaviors that make them less suitable for our tasks. We encourage the reader to visit the project webpage for videos of our real-world rollouts.
Figure 7: Failure Analysis. We report the failure case analysis for each method, aggregated across all real-world experiments.
Failure Analysis. We analyze the failure modes of the methods evaluated and present the results in Figure 7. Failures are categorized as: runs aborted automatically or manually for hardware safety (“abort”), lack of motion (“stuck”), incorrect object localization (“coarse misgrasp”), or failed manipulation (“fine misgrasp”), alongside other cases (“others”). Notably, MAPLE is the only method that never caused an abort and showed the least localization errors (counting coarse and fine misgrasps). The analysis confirms the superior efficacy and safety of our encoder against baselines for dexterous manipulation tasks in the real world.
@@ 5 Conclusion
In this work, we studied how to infuse dexterous manipulation priors into visual representations by training on diverse human videos, aiming to improve performance on downstream robotic tasks in human-centric environments. We introduced MAPLE, a novel approach that learns manipulation-specific priors from egocentric videos. MAPLE trains an encoder to predict hand-object contact points and detailed 3D hand poses from a single image, producing features useful for downstream dexterous manipulation tasks. For the benefit of the community and a more rigorous evaluation, we further proposed new dexterous simulation environments. Our strong results on both new tasks and existing benchmark tasks in simulation, as well as our superior real-world performance demonstrated promising manipulation prior learning and excellent generalization ability with MAPLE. Future directions of research include combining MAPLE with language conditioning and incorporating more diverse, possibly exocentric human manipulation datasets during its training.
@@ 6 Acknowledgements
This work was supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a03 on Alps, as well as the Swiss National Science Foundation Advanced Grant 216260: “Beyond Frozen Worlds: Capturing Functional 3D Digital Twins from the Real World”.
[…]
@@
Figure S8: The initial setups of all real-world evaluations: open the bottle cap (bottle), place the pan (pan), and wash the dish (sponge). Each method is evaluated for 16 episodes, starting from each of the states shown in the figure.
In this section, we provide more details on the imitation learning evaluations introduced in Sec. 4.3. Each method is evaluated for 16 episodes for every task, with predefined initial setups as visualized in Fig. S8.
@@
Our system consists of an ORCA robotic hand [ORCA] mounted on a Franka Emika Panda robotic arm [FrankaEmika]. One external and one wrist-mounted OAK-D camera [OakD] capture the scene, and the arm and hand are controlled at 10 Hz. The arm’s action is defined by a 7-dimensional end-effector pose (x, y, z, plus a quaternion), and the hand’s action is represented by a 17-dimensional vector corresponding to its joint positions. All camera images are cropped and resized to $480\times 270$ pixels. Our framework uses an observation window spanning two time steps, while control actions are executed in chunks spanning 10 time steps.
S1.2 Failure Cases
Coarse/Fine Misgrasp. In Fig. 7, we define “coarse misgrasp” as an episode in which the policy moves the robot arm toward a direction not leading to the target object, closes the hand without grasping anything, then either stops moving the arm in a significant manner for over 15 seconds or fails to touch the object at least 6 times after attempting to grasp the object again. In case the policy moves the arm within the hand’s grasping range, or the hand touches the object but fails to lift it, a “fine misgrasp” classification follows instead.
Stuck. An episode in which the hand never closes its grasp yet the arm stops significant movement for at least 15 seconds is classified as “stuck”.
Abort. Additionally, the policy execution framework automatically aborts the episode in case the arm moves outside the defined joint limits, a maximum joint velocity is exceeded, or the difference between the measured and expected joint position grows too large (e.g. as a result of the arm colliding with an obstacle). Alternatively, the human operator may stop the execution manually if they detect a dangerous deformation of the hand’s fingers due to an incorrect interaction with an object, or in case of an imminent collision of the hand with a wall. In such an event, the episode is classified as “abort”.
Other. Rarely observed reasons for failure classified as “other” may include knocking over the object, placing it in an unintended target location, inability to open the cap of a bottle despite grasping it due to not strong enough movements, or similar.
S1.3 Imitation Learning Details
To accelerate training and enable larger batch sizes, we cache embeddings from the observation encoders directly within the dataset. This approach significantly speeds up training. To increase the policy’s robustness to visual perturbations, we create 10 cached copies for each observation, each augmented differently.
@@
Our DiT [dasari2024DiTpolicy] backbone has a model dimension of 512, consists of 6 blocks, and employs 8 attention heads. The network is trained using the Adam [Adam] optimizer with a cosine warmup learning rate schedule. All policies working in the real world are evaluated after 35,000 gradient steps.
S1.4 Detailed Task Descriptions
We report details regarding the demonstration collection and evaluation of each of our three real-world tasks.
S1.4.1 “Open the Bottle Cap” Task
@@
Demonstrations. 285 demonstrations are collected and used for training. 80% of them are used as training set, 20% as evaluation set. The starting position of the bottle is randomized. Demonstrations are collected by an experienced operator wearing Rokoko Smartgloves and the Coil Pro motion capture setup [Smartglove], which provides absolute proprioceptive data of the demonstrator’s fingers and wrist positions. Images from two cameras are collected and fed during policy rollout: front camera and wrist camera.
Evaluation. Each method is evaluated for 16 episodes, all starting from the same set of initial conditions as shown in Fig. S8.a. Before each trial, the arm and hand are moved to a normative pose. An episode is counted as successful only if the bottle cap is opened using the fingers. Due to the difficulty of the task, the policy is allowed to reattempt opening the cap up to 6 times if it misses the bottle.
S1.4.2 “Place the Pan” Task
Demonstrations. 450 demonstrations are collected and used for training. 80% of them are used as training set, 20% as evaluation set. The position and orientation of the pan and stove on the table is randomized. The demonstrations are collected with the same Rokoko solution as described before. Images from two cameras are collected and fed during policy rollout: front camera and wrist camera.
Evaluation. Each method is evaluated for 16 episodes, all starting from the same set of initial conditions as shown in Fig. S8.b. Before each trial, the arm and hand are moved to a normative pose. An episode is counted as successful only if the pan is grasped, lifted and placed stably on the induction stove.
S1.4.3 “Wash the Dish” Task
Demonstrations. 227 demonstrations are collected and used for training. 80% of them are used as training set, 20% as evaluation set. The starting position of the sponge and plate together is randomized. The demonstrations are collected using the same Rokoko solution as described above. Images from two cameras are collected and fed during the policy rollout: front camera and wrist camera.
Evaluation. Each method is evaluated for 16 episodes, all starting from the same set of initial conditions as shown in Fig. S8.c. Before each trial, the arm and hand are moved to a normative pose. An episode is counted as successful only if the sponge is grasped, lifted, placed on the plate and rotated around the plate’s interior at least three times.
S2 Zero-Shot Real World Experiments
To test the robustness of policies trained using our features against those using features of baseline methods, we devise two zero-shot experiments in which the policies are exposed to out-of-domain visual inputs without corresponding training data and measure the resulting success rates. Specifically, we evaluate the methods on a version of the “Wash the Dish” task featuring a sponge with a different color and shape. Moreover, we add an additional pan to the “Place the Pan” task and evaluate the policies. We conduct 16 trials for each method and task, following the positions visualized in Figure S8. Exemplary configurations are visualized in Figure S9.
We summarize the result in Table S4. The policies using our features outperform those using competing encoders, demonstrating the promising zero-shot capabilities and robustness of MAPLE.
Figure S9: Visualization of Zero-Shot Initial Configurations. Exemplary initial configurations for our zero-shot task variations. We modify the “Wash the Dish” task to use a sponge of a different color and shape than seen during training (“Unseen Object”). We further conduct evaluations on the “Place the Pan” task when an additional pan is present in the image as a “Distractor Object”.
Experiment DINO HRP VC-1 R3M MAPLE
Unseen Object 25.0% 0% 0% 25.0% 31.3%
Distractor Object 0% 0% 0% 0% 18.8%
Table S4: Performance on Zero-Shot Experiments. We report success rates for MAPLE and four baseline methods on our zero-shot task variations. See Figure S9 for visualizations of some initial configurations.
S3 Demonstration Data and Benchmark Environments for Simulation
@@
Figure S11: Example Demonstration Trajectory: Camera observations of a demonstration for the ’Drill’ task. The top row displays the left camera view, the middle row shows the right camera view, and the bottom row presents the top camera view.
@@
S3.1 Hand Retargeting Algorithm
@@
The algorithm from [DexMV] works by optimizing over the vector of the Adroit joint actuations, minimizing a Huber loss on the difference between pairs of an inter-fingertip or finger-to-wrist vector on the human hand and a corresponding vector measuring the same distance on the robotic hand. The hand pose is inferred by the aforementioned retargeting procedure, while the global position of the wrist joint is obtained using the Rokoko Coil Pro, an EMF-based tracking device. We use virtual PID controllers to drive each joint.
S3.2 Visualization of Demonstration Trajectories
@@
We show selected frames from our demonstration trajectories for the brush (Figure S10), the drill (Figure S11), the iron (Figure S12) and the pan (Figure S13) tasks for each camera view.
S4 Evaluation Details in Simulation
@@
To evaluate the suitability of a given visual encoder for downstream dexterous manipulation tasks, we use a simulator in which a policy network consuming features from a visual encoder is trained to operate an observed robotic manipulator. As usual in a real-world deployment, the policy network additionally receives proprioceptive features describing the current joint configurations of the robot as input. Position and orientation information related to the object(s) to be manipulated is only provided through visual cues, so as to increase realism and encourage the use of visual features to grasp and manipulate the object. All considered environments feature the Adroit [AdroitHand] dexterous hand with 30 DoF in total, which is allowed to move freely within a predefined operating range above the table holding the object(s) to manipulate.
@@
The policy networks are trained to perform the respective tasks using demonstrations recorded by human participants, as described in Sec. 4.2. While training the policy networks, we periodically perform rollouts (i.e. simulation runs where the performance of the policy is assessed). The policies are trained for 20,000 training iterations. Rollouts are performed every 1,000 training iterations, and the highest success rate achieved over a given run is chosen for the computation of the final average success rates reported in the tables. 50 rollouts are used during each evaluation, and we calculate the success rate for a given policy and task by averaging over the success rates obtained using 3 different viewpoints and 3 policy seeds per viewpoint. We randomize the position and orientation of the object(s) to manipulate upon each evaluation start to make the tasks more challenging and realistic.
@@
During the training of the policy, the checkpoint of the respective visual encoder is used as a frozen feature extractor, and only the parameters of the policy are trained. For the MAPLE model and all ablations, we use the checkpoint obtained after training 3,000 iterations on our Ego4D-based dataset. This iteration count was chosen to match the approximate contact prediction performance saturation point on a small validation dataset, manually constructed from prediction frames produced by our data extraction pipeline. In all experiments, the policy is a 2-layer MLP with 256 units for each hidden layer. The MLP receives the visual features produced by the respective encoder from a rendered RGB view of the input scene. Examples of rendered simulator views are visualized in Figure 3. The MLP further receives a number $p$ of proprioceptive features encoding the current joint configuration of the jaw gripper resp. robotic hand. Specifically, due to the hand being fixed, we use $p=24$ for the DAPG pen environment, and $p=30$ for all other environments due to the hand’s movement adding 6 additional DoF. All these values are chosen to match the setup in previous work [R3M]. The number of proprioceptive features generally corresponds to the DoF of the robotic manipulator. At each simulation timestep, given these environment observations, the MLP networks output actuation vectors which determine the force applied at each actuator according to the conversion outlined in [MujocoXMLReference]. To learn from human demonstrations, the policy is provided observation inputs recorded during human teleoperation of the robot manipulator, and is trained to predict the actuation vector corresponding to the action that the human performed at that timestep. As is standard in literature [DAPG, R3M], we train a separate policy per environment and use the same base seed for all object position/orientation randomizations to ensure reproducibility.
S4.1 Environment Success Criteria
@@
Here, we provide detailed criteria that must be fulfilled for a task to count as successfully completed, for each of our dexterous environments. For distance measurements, MuJoCo site markers [MujocoXMLReference] attached to the robotic hand, manipulated object (brush, drill, iron, pan) and target object (wooden board, clothes, induction stove) are used.
  • Drill: The index finger must be in close vicinity of the drill’s trigger, and the tip of the drill bit must be in close vicinity of the wooden board’s front side.
  • Iron: The bottom of the clothes iron must slide along at least three of eight markers placed equidistantly along the right leg of the pants laid out on the table, while the palm of the hand is in close vicinity of the handle of the clothes iron.
@@
S5 Contact Point Prediction Details
Here, we provide definitions for the SIM and NSS metrics used in Sec. 4.1, training details, as well as qualitative examples of predictions obtained using our method and a representative baseline.
S5.1 Metric Definitions
Following the evaluation protocol of [liu2022joint], for Tab. 1 we train a conditional variational autoencoder [cVAE] working with features from various encoders and predicting $c=5$ contact points per given input image showing a person about to interact with some object of a scene. Setting $c>1$ helps account for the ambiguity of this contact point prediction, arising from multiple possible actions. To evaluate the performance of the resulting contact point predictor, each predicted contact point for the given image is rendered as an isotropic Gaussian with $\sigma=3$ on a heatmap $M\in\mathbb{R}^{H\times W}$ of height $H$ and width $M$. The ground-truth contact point is rendered in an identical manner on a ground-truth heatmap $\hat{M}\in\mathbb{R}^{H\times W}$. Given $M$ and $\hat{M}$, we define $SIM$ and $NSS$ for the corresponding image as follows:
[…]
$NSS(M,\hat{M})=\sum_{y=1}^{H}\sum_{x=1}^{W}\mathbb{I}({\hat{M}_{y,x}>0})\cdot\frac{M_{y,x}-\mu_{M}}{\sigma_{M}},$
where
$n_{K}=\sum_{y=1}^{H}\sum_{x=1}^{W}K_{y,x},\ \mu_{K}=\frac{n_{K}}{H\times W},$
$\sigma_{K}=\sqrt{\frac{\sum_{y=1}^{H}\sum_{x=1}^{W}(K_{y,x}-\mu_{K})^{2}}{H\times W}},$
and $\mathbb{I}$ is the indicator function. The final $SIM$ and $NSS$ metrics are averages over their image-wise values for the dataset used in [liu2022joint]. Following the official implementation of [liu2022joint], we set H=32,W=32H=32,\ W=32.
S5.2 Training Details
We train the cVAE contact predictors for 3000 iterations, evaluating them every 150 iterations. A method’s score in Tab. 1 is determined based on the checkpoint that reached the highest SIM score during all evaluations of the method. The losses used are identical to those in [liu2022joint]. We use AdamW [AdamW] with a learning rate of $5\times 10^{-5}$ and a batch size of 128.
S5.3 Qualitative Examples
We visualize predictions for two samples of our contact point predictor trained using MAPLE features vis-à-vis those of the DINO-based contact predictor in Figure S14. Here, we only incorporate the top-scoring prediction in the heatmap for visual clarity. We further show additional high-quality predictions by the MAPLE-based contact predictor in Figure S15.
(a) DINO
(b) MAPLE
(c) Ground-Truth
(d) DINO
(e) MAPLE
(f) Ground-Truth
Figure S14: Qualitative Comparison of Contact Predictions. We showcase the output of cVAE contact predictor heads working with DINO vs. MAPLE features for two samples from [liu2022joint].
Figure S15: Further Qualitative Examples of Contact Predictions. Our encoder allows for high-quality contact point predictions on egocentric videos when combined with a cVAE predictor head.
S6 Training Supervision Extraction
@@
To obtain data for training our visual encoder, we start by processing every frame in the clips subset of the Ego4D dataset [Ego4D] using the VISOR-HOS [EKVisor] hand-object interaction segmenter, and retain frames where exactly one right hand is contacting an object with an HOS confidence at least $c\geq 0.9$, and no left hand is contacting an object with a similarly high confidence. Afterwards, we obtain frame-wise HaMeR [HaMeR] hand reconstructions for each contact frame. These hand reconstructions are used to initialize the prediction-frame–seeking contact point tracking with the thumb and index fingertips: for each contact frame, we first perform a binary erosion operation (12 iterations) on the object mask to remove the boundary, as motivated in Section 3.2. Then, we project the thumb and index fingertip points of the right hand in the contact frame to the eroded object mask. In case the ratio of the projected points’ Euclidean distance to their original distance falls outside the range between 0.3 and 1.7, the contact frame is not used. This helps eliminate objects with degenerate HOS masks. Otherwise, we initialize SAM-PT [SamPT] with the projected fingertip points, as well as 10 other points sampled from the eroded object mask according to the query point sampling algorithm described in [SamPT] and used to track the object. We continue tracking the points backward through the video until we encounter a frame where the HOS hand mask dilated using 75 binary dilation iterations (i.e., the expanded hand mask) no longer contains the tracked projected fingertip points. In case no such frame is encountered after 45 frames (i.e. 1.5s), the given contact frame is discarded. Otherwise, we use this frame as the prediction frame, and the backtracked points in that frame as training supervision for the contact loss. For the hand pose loss, we note that there are little constraints on the hand pose as it is moving in the direction of an object to grasp it. Hence, we always let our method reconstruct the hand pose at the contact frame, to promote learning the more informative grasping pose expected to be seen when the hand is already in contact with the object.
@@
Figure S17: Visualization of Object Mask and Contact Points. The left column displays the prediction frames, while the right column shows the corresponding contact frames automatically extracted during our dataset acquisition step. The cyan cross mark represents the contact point of the thumb, and the pink cross mark indicates the contact point of the index finger.
S6.1 Error Analysis of Contact Supervision Extraction